# RLHF Workflow: From Reward Modeling to Online RLHF

### v2

@@ Abstract
We present the workflow of Online Iterative Reinforcement Learning from Human Feedback (RLHF) in this technical report, which is widely reported to outperform its offline counterpart by a large margin in the recent large language model (LLM) literature. However, existing open-source RLHF projects are still largely confined to the offline learning setting. In this technical report, we aim to fill in this gap and provide a detailed recipe that is easy to reproduce for online iterative RLHF. In particular, since online human feedback is usually infeasible for open-source communities with limited resources, we start by constructing preference models using a diverse set of open-source datasets and use the constructed proxy preference model to approximate human feedback. Then, we discuss the theoretical insights and algorithmic principles behind online iterative RLHF, followed by a detailed practical implementation. Our trained LLM, LLaMA-3-8B-SFR-Iterative-DPO-R, achieves impressive performance on LLM chatbot benchmarks, including AlpacaEval-2, Arena-Hard, and MT-Bench, as well as other academic benchmarks such as HumanEval and TruthfulQA. We have shown that supervised fine-tuning (SFT) and iterative RLHF can obtain state-of-the-art performance with fully open-source datasets. Further, we have made our models, curated datasets, and comprehensive step-by-step code guidebooks publicly available. Please refer to https://github.com/RLHFlow/RLHF-Reward-Modeling and https://github.com/RLHFlow/Online-RLHF for more detailed information.
@@ 1 Introduction
Reinforcement Learning from Human Feedback (RLHF) has become as a key technique for integrating human preference signals into machine learning methods, particularly in aligning Large Language Models (LLMs) with human values and preferences (Christiano et al.,, 2017; Ziegler et al.,, 2019). Notable examples include the revolutionary closed-source ChatGPT (OpenAI,, 2023), Claude (Anthropic,, 2023), and Gemini (Team et al.,, 2023), as well as the powerful open-source models like Zephyr (Tunstall et al.,, 2023), Starling (Zhu et al.,, 2023), and LLaMA-3 (Meta,, 2024). In particular, since the introduction of ChatGPT, RLHF has attracted significant interest in a diverse set of communities. However, compared to the supervised fine-tuning that is rather well studied with many great open-source projects like Open-Hermes (Teknium,, 2023) and Vicuna (Zheng et al.,, 2023), RLHF remains relatively under-explored within the open-source community.
To facilitate our discussion, we build upon the standard RLHF workflow (Ouyang et al.,, 2022; Bai et al., 2022b, ; Touvron et al.,, 2023). We characterize an LLM by a policy $\pi$, which takes a prompt $x\in\mathcal{X}$ and produces a response $a\in\mathcal{A}$ from the distribution $\pi(\cdot|x)$. We denote the initial model of RLHF as $\pi_{0}$, which is fine-tuned on some instruction-following data after the pre-training stage. We assume that we have a prompt set that is sampled from some unknown but fixed distribution $x\sim d_{0}$. The key component in RLHF is the Preference Oracle, which is mathematically defined as follows.
@@
To further simplify the problem, it is commonly assumed that the preference signal can be modeled using the reward-based Bradley-Terry model, a well-known approach in preference learning (Bradley and Terry,, 1952; Ouyang et al.,, 2022; Bai et al., 2022a, ; Touvron et al.,, 2023).
@@
This modeling is a proxy of preference oracle and connects the learning objective of RLHF with reward maximization. In practice, since the BT model may not fully capture the complex human preference, we usually optimize the following KL-regularized target:
@@
where $\eta>0$ is the KL penalty coefficient. This formulation is widely studied in practice (Ziegler et al.,, 2019; Wu et al.,, 2021; Ouyang et al.,, 2022; Rafailov et al.,, 2023; Liu et al., 2023a, ; Xiong et al.,, 2023) and admits the following intractable closed-form solution (Zhang,, 2023)
@@ 1.1 Previous RLHF Algorithms and Their Challenges
Broadly speaking, previous RLHF methods can be largely divided into two categories: (1) deep RL-based approach using Proximal Policy Optimization (PPO) (Schulman et al.,, 2017; Christiano et al.,, 2017; Ziegler et al.,, 2019) and (2) (offline) direct preference learning (e.g., DPO) approaches (Zhao et al.,, 2023; Rafailov et al.,, 2023; Azar et al.,, 2023; Tang et al.,, 2024).
DRL-based framework. The DRL-based framework consists of two stages. In the first stage, a reward model is trained. Specifically, given a preference dataset $\mathcal{D}_{\mathrm{off}}=\{(x,a^{w},a^{l})\}$, where $a^{w}$ is a response preferred over $a^{l}$ given the instruction or prompt $x$. The log-likelihood function of the BT model can be expressed as follows:
@@
We can compute the maximum likelihood estimator (MLE) $r_{\mathrm{MLE}}$ based on $\mathcal{D}_{\mathrm{off}}$ by maximizing the $\ell_{\mathcal{D}_{\mathrm{off}}}(\theta)$. Then, in the second stage, DRL methods like PPO can be applied to optimize against the following regularized reward:
$\hat{r}(x,a)=r_{\mathrm{MLE}}(x,a)-\eta\log\frac{\pi(a|x)}{\pi_{0}(a|x)}.$
@@
This approach has been employed by ChatGPT (Ouyang et al.,, 2022) and Claude (Bai et al., 2022a, ) and has contributed to the alignment of LLaMA-2/3 (Touvron et al.,, 2023). However, it is known that even in the best case, tuning the DRL method to its best performance requires extensive efforts in hyper-parameter selection and code-level optimization (Choshen et al.,, 2019; Engstrom et al.,, 2020). This becomes even more challenging in the context of LLMs, as fine-tuning LLMs is computationally expensive. Additionally, the PPO algorithm requires loading multiple LLMs simultaneously, including the actor (policy), critic (value network), reward model, and reference model (for KL estimation), which places significant pressure on GPU memory, especially for resource-constrained open-source projects.
@@
Direct preference learning. In view of the above issues of PPO, there is an innovative line of work that directly learns from human preference datasets without explicitly constructing a reward function (Zhao et al.,, 2023; Rafailov et al.,, 2023; Azar et al.,, 2023). Among these methods, the direct preference optimization (DPO) algorithm is particularly popular. It leverages Equation 3 to formulate reward as a function of policy and directly optimizes the following loss function using the preference dataset $\mathcal{D}_{\mathrm{off}}$:
$\mathcal{L}_{\mathcal{D}_{\mathrm{off}}}(\theta,\pi_{0})=-\sum_{(x,a^{w},a^{l})\in\mathcal{D}_{\mathrm{off}}}\Big{[}\log\sigma\Big{(}\eta\log\frac{\pi_{\theta}(a^{w}|x)}{\pi_{0}(a^{w}|x)}-\eta\log\frac{\pi_{\theta}(a^{l}|x)}{\pi_{0}(a^{l}|x)}\Big{)}\Big{]}.$ (5)
@@
Direct preference learning algorithms are generally easier to tune and require fewer computational resources compared to DRL methods. Notably, in many LLM benchmarks’ leaderboards (Dubois et al.,, 2023; Zheng et al.,, 2023), powerful open-source models are primarily aligned using DPO. Considering these factors, in our project, we focus on direct preference learning algorithms while leaving the study of the DRL-based framework for future research.
@@
While the vanilla offline direct preference learning algorithms are useful in some case studies, they also face certain challenges. Specifically, they are considered offline because they learn from an offline preference dataset collected through the following process:
$x\sim d_{0},a^{1}\sim\pi_{D}^{1},a^{2}\sim\pi_{D}^{2},\qquad y\sim\mathrm{Ber}\big{(}\mathbb{P}(a^{1}\succ a^{2}|x,a^{1},a^{2})\big{)}.$ (6)
@@
Here, $(\pi_{D}^{1},\pi_{D}^{2})$ represent two behavior policies, often taken as $\pi_{0}$, other open-sourced models or proprietary models. The term “offline learning” implies that we cannot further query the preference oracle $\mathbb{P}$ during the training process. However, the finite dataset $\mathcal{D}_{\mathrm{off}}$ used for offline learning leads to the problem of over-optimization. This is because the finite dataset fails to cover the entire prompt-response space and the resulting policy model often performs poorly when faced with out-of-distribution data (Burns et al.,, 2023). Indeed, Xiong et al., (2023) shows that we need a uniform coverage of the space to make sure that the vanilla DPO (Rafailov et al.,, 2023) can learn the optimal policy. Such a coverage condition is hard to satisfy in practice due to the exponentially large space. Here are a few statistics that highlight this issue.
  • Along the way of the RLHF training, the average density ratio $\frac{\pi(a|x)}{\pi_{0}(a|x)}>\exp(25)$ as reported in Figure 13 of Bai et al., 2022a . See similar results of rejection sampling fine-tuning (Dong et al.,, 2023) and DPO (Rafailov et al.,, 2023).
  • As a case study, consider using the behavior policy Gemma-7B-it to collect data for aligning Gemma-2B-it (Team et al.,, 2024) using 15k prompts from (Cui et al.,, 2023). The average KL divergence between Gemma-7B-it and Gemma-2B-it is calculated to be $456.4$.
Therefore, the distribution shift between policies is usually very large, and it is unlikely that we can learn the optimal policy solely from a pre-collected dataset.
@@ 1.2 Online Iterative RLHF
In contrast, the Claude project (Bai et al., 2022a, ) and LLaMA-2 project (Touvron et al.,, 2023) have demonstrated that online iterative RLHF can significantly improve model performance. The process of online iterative RLHF, as formally formulated in Xiong et al., (2023), can be summarized as follows. Given the pre-collected preference dataset $\mathcal{D}=\mathcal{D}_{\mathrm{off}}$ (if applicable, otherwise empty), for each iteration $t\in[T]$:
@@
The effectiveness of online data can be intuitively explained in the context of RLHF as follows. Initially, most of the samples from policy $\pi_{0}$ are in a low-reward regime and are used for the reward modeling. As a result, our reward model is accurate in evaluating low-reward responses. As training progresses, however, the policy improves and starts to generate responses with higher rewards. These high-reward responses may fall into the out-of-distribution (OOD) domain of the reward function, where the constructed reward model is no longer reliable. In contrast, using the intermediate policy to sample new responses, querying human feedback on these samples, and add them back to the training set can significantly mitigate the OOD problem, which traditional offline methods may not handle effectively. This approach enhances the reliability of the reward model in the high-reward regime, which in turn improves the policy’s performance. The intuition extends to the direct preference learning algorithms, where we essentially enlarge our preference dataset by continuously collecting new online data.
The great work of Huggingface (Tunstall et al.,, 2023) provides an open-source recipe to do direct preference learning from high-quality offline preference dataset, which can be viewed as an efficient way of distillation. In contrast, the online RLHF is still largely under-explored in the literature. Xiong et al., (2023) made the first step towards understanding the advantage of online exploration in RLHF from a theoretical perspective, and this technical report aims to provide a detailed recipe to verify the effectiveness of the proposed framework.
@@ 1.3 Human Feedback Approximation
Ideally, the online preference signal is sampled from a representative group of human labelers. However, human feedback is extremely expensive in practice, which the open-source community usually cannot afford. In the literature, there is a line of work showing that training a proxy preference model, and using the preference model to give proxy labels in a semi-supervised manner improve the model performance (Dong et al.,, 2023; Yuan et al.,, 2023; Liu et al., 2023a, ; Hoang Tran,, 2024). We conjecture that this is because the reward model (discriminator) usually generalizes better than the policy (generator).
In particular, Hoang Tran, (2024) shows that if the preference model (reward model) is trained on a diverse set of preference datasets, the Pair-RM (Jiang et al.,, 2023) with only 0.4B parameters can provide iterative preference learning with meaningful signals so that the resulting model [^1] https://huggingface.co/snorkelai/Snorkel-Mistral-PairRM-DPO achieves an impressive AlpacaEval-2 length-control win rate of 26.4%. Motivated by this line of work, we first train a proxy preference (reward) model based on the diverse open-source preference datasets in Section 2 and then use the resulting model to provide preference signals for the subsequent iterative RLHF.
@@ 1.4 Related Work
RLHF and RLHF algorithms. The dominant RLHF framework used for the LLM alignment was first popularized in Christiano et al., (2017); Ziegler et al., (2019) and was further developed in Instruct-GPT (Ouyang et al.,, 2022), Claude (Bai et al., 2022a, ), and LLaMA-2 (Touvron et al.,, 2023). These works typically involve constructing a reward model based on the MLE of the Bradley-Terry model, and then using the PPO algorithm to optimize the reward signals with KL regularization. One notable exception is that the LLaMA-2 uses a mixture of rejection sampling fine-tuning (Dong et al.,, 2023; Wang et al.,, 2024) and PPO in their RLHF pipeline. We refer the interested readers to Bai et al., 2022a ; Touvron et al., (2023) for a detailed description. However, the use of PPO in RLHF has limitations. It is known to be unstable (Choshen et al.,, 2019), sensitive to implementation (Engstrom et al.,, 2020), and resource-intensive (Yuan et al.,, 2023). Despite some efforts to improve PPO in the context of RLHF (Li et al.,, 2023; Chan et al.,, 2024; Chang et al.,, 2024; Zhong et al.,, 2024), reproducing the successful results achieved with PPO is challenging for the open-source community due to these limitations as it requires extensive efforts and resources that the open-source communities usually cannot afford. In recognition of these issues of PPO, a line of work studies the (offline) direct preference learning algorithms, including Slic (Zhao et al.,, 2023), DPO (Rafailov et al.,, 2023), IPO (Azar et al.,, 2023), KTO (Ethayarajh et al.,, 2024), ARM (Pang et al.,, 2024), and GPO (Tang et al.,, 2024). These algorithms skip the reward modeling step, and optimize a designed loss target on the offline preference dataset directly (hence the name). It is widely observed that the direct preference learning algorithms are much more stable than the PPO, and achieve impressive performance evaluated by standard benchmarks (Tunstall et al.,, 2023; Dubois et al.,, 2023; Zheng et al.,, 2023).
@@
RLHF benefits from online (iterative) learning. Roughly speaking, online iterative learning means that we will deploy the intermediate models and query human feedback for the responses of these models. Intuitively, this strategy can help to mitigate the OOD issue of the learned reward model (Gao et al.,, 2023), and its advantages have been reported in Ouyang et al., (2022); Touvron et al., (2023) for the PPO-based framework. Even when the additional feedback is derived from a proxy reward constructed from the same offline dataset (similar to semi-supervised learning), iterative rejection sampling fine-tuning (RAFT) (Dong et al.,, 2023) and DPO based on samples from the target distribution estimator have been shown to outperform the original offline counterparts (Pang et al.,, 2024; Liu et al., 2023a, ). Furthermore, recent works (Xiong et al.,, 2023; Xu et al.,, 2023; Hoang Tran,, 2024; Yuan et al., 2024b, ; Swamy et al.,, 2024; Chen et al., 2024b, ; Ye et al.,, 2024; Guo et al.,, 2024; Rosset et al.,, 2024; Tajwar et al.,, 2024; Calandriello et al.,, 2024; Wu et al.,, 2024) have demonstrated that online iterative variants of direct preference learning algorithms significantly outperform their offline counterparts. In particular, we refer the interested readers to Guo et al., (2024) for the extensive experimental results with different offline base algorithms.

### v3

@@ Abstract
We present the workflow of Online Iterative Reinforcement Learning from Human Feedback (RLHF) in this technical report, which is widely reported to outperform its offline counterpart by a large margin in the recent large language model (LLM) literature. However, existing open-source RLHF projects are still largely confined to the offline learning setting. In this technical report, we aim to fill in this gap and provide a detailed recipe that is easy to reproduce for online iterative RLHF. In particular, since online human feedback is usually infeasible for open-source communities with limited resources, we start by constructing preference models using a diverse set of open-source datasets and use the constructed proxy preference model to approximate human feedback. Then, we discuss the theoretical insights and algorithmic principles behind online iterative RLHF, followed by a detailed practical implementation. Our trained LLM achieves impressive performance on LLM chatbot benchmarks, including AlpacaEval-2, Arena-Hard, and MT-Bench, as well as other academic benchmarks such as HumanEval and TruthfulQA. We have shown that supervised fine-tuning (SFT) and iterative RLHF can obtain state-of-the-art performance with fully open-source datasets. Further, we have made our models, curated datasets, and comprehensive step-by-step code guidebooks publicly available.
@@ 1 Introduction
Reinforcement Learning from Human Feedback (RLHF) (Christiano et al., 2017; Ziegler et al., 2019) has become as a key technique for integrating human preference signals into machine learning methods. In particular, RLHF has become a standard component in the post-training pipe line of foundation Large Language Models (LLMs), which serves to align the outputs of these models with human values such as helpfulness, harmlessness, and honesty (Ouyang et al., 2022; Bai et al., 2022a). Notable examples include the revolutionary closed-source ChatGPT (OpenAI, 2023), Claude (Anthropic, 2023), and Gemini (Team et al., 2023), as well as the powerful open-source models like Zephyr (Tunstall et al., 2023), Starling (Zhu et al., 2023), and LLaMA-3 (Meta, 2024). In particular, since the introduction of ChatGPT, RLHF has attracted significant interest in a diverse set of communities. However, compared to the supervised fine-tuning that is rather well studied with many great open-source projects like Open-Hermes (Teknium, 2023) and Vicuna (Zheng et al., 2023), RLHF remains relatively under-explored within the open-source community.
To facilitate our discussion, we build upon the standard RLHF workflow (Ouyang et al., 2022; Bai et al., 2022b; Touvron et al., 2023). We characterize an LLM by a policy $\pi$, which takes a prompt $x\in\mathcal{X}$ and produces a response $a\in\mathcal{A}$ from the distribution $\pi(\cdot|x)$. We denote the initial model of RLHF as $\pi_{0}$, which is fine-tuned on some instruction-following data after the pre-training stage. We assume that we have a prompt set that is sampled from some unknown but fixed distribution $x\sim d_{0}$.
The central principle of RLHF is to learn from relative feedback, rather than an absolute reward signal, as is common in traditional RL literature. This approach is preferred because human raters often struggle to provide accurate absolute ratings; instead, they find it easier to choose between two options, indicating which response they prefer (Christiano et al., 2017). This idea can further date back to the study of dueling bandit (Joachims et al., 2007; Yue et al., 2012) in the context of online decision making. Specifically, we assume that we have access to a Preference Oracle, that is a proxy of a real-world human rater. Mathematically, we have the following formal definition.
@@
To further simplify the problem, it is commonly assumed that the preference signal can be modeled using the reward-based Bradley-Terry model, a well-known approach in preference learning (Bradley & Terry, 1952; Ouyang et al., 2022; Bai et al., 2022a; Touvron et al., 2023).
@@
Although the BT model may not fully capture the complex human preference, it tends out to be a useful approximation to connect the learning objective of RLHF with reward maximization and has achieved tremendous success in making ChatGPT (Ouyang et al., 2022) and Claude (Bai et al., 2022a). Meanwhile, in response to the imperfect nature of BT model, the goal in this RLHF formulation is to optimize the following KL-regularized target:
@@
where $\eta>0$ is the KL penalty coefficient. This formulation is widely studied in practice (Ziegler et al., 2019; Wu et al., 2021; Ouyang et al., 2022; Rafailov et al., 2023; Liu et al., 2023a; Xiong et al., 2023) and admits the following intractable closed-form solution (Zhang, 2023)
@@ 1.1 Previous RLHF Algorithms and Their Challenges
Generally, previous RLHF methods can be largely divided into two categories: (1) deep RL-based approach using Proximal Policy Optimization (PPO) (Schulman et al., 2017; Christiano et al., 2017; Ziegler et al., 2019) and (2) (offline) direct preference learning (e.g., DPO) approaches (Zhao et al., 2023; Rafailov et al., 2023; Azar et al., 2023; Tang et al., 2024).
DRL-based framework. The DRL-based framework consists of two stages. In the first stage, a reward model is trained. Specifically, given a preference dataset $\mathcal{D}_{\mathrm{off}}=\{(x,a^{w},a^{l})\}$, where $a^{w}$ is a response preferred over $a^{l}$ given the instruction or prompt $x$. The log-likelihood function of the BT model can be expressed as:
@@
We can compute the maximum likelihood estimator (MLE) $r_{\mathrm{MLE}}$ based on $\mathcal{D}_{\mathrm{off}}$ by maximizing the $\ell_{\mathcal{D}_{\mathrm{off}}}(\theta)$. In the second stage, DRL methods like PPO can be applied to optimize against the following reward:
$\small\hat{r}(x,a)=r_{\mathrm{MLE}}(x,a)-\eta\log\frac{\pi(a|x)}{\pi_{0}(a|x)}.$
@@
This approach has been employed by ChatGPT (Ouyang et al., 2022) and Claude (Bai et al., 2022a) and has contributed to the alignment of LLaMA-2 (Touvron et al., 2023). However, it is known that even in the best case, tuning the DRL method to its best performance requires extensive efforts in hyper-parameter selection and code-level optimization (Choshen et al., 2019; Engstrom et al., 2020). This becomes even more challenging in the context of LLMs, as fine-tuning LLMs is computationally expensive and searching the complicated hyper-parameters configuration is generally infeasible. Additionally, the PPO algorithm requires loading multiple LLMs simultaneously, including the actor (policy), critic (value network), reward model, and reference model (for KL estimation), which places significant pressure on GPU memory, especially for resource-constrained open-source projects.
@@
Direct preference learning. In view of the above issues of PPO, there is an innovative line of work that directly learns from human preference datasets without explicitly constructing a reward function (Zhao et al., 2023; Rafailov et al., 2023; Azar et al., 2023). Among these methods, the direct preference optimization (DPO) algorithm is particularly popular. It leverages Equation 3 to formulate reward as a function of policy and directly optimizes the following loss function using the preference dataset $\mathcal{D}_{\mathrm{off}}$:
$\small\mathcal{L}_{\mathcal{D}_{\mathrm{off}}}(\theta,\pi_{0})=-\sum_{(x,a^{w},a^{l})\in\mathcal{D}_{\mathrm{off}}}\Big{[}\log\sigma\Big{(}\eta\log\frac{\pi_{\theta}(a^{w}|x)}{\pi_{0}(a^{w}|x)}-\eta\log\frac{\pi_{\theta}(a^{l}|x)}{\pi_{0}(a^{l}|x)}\Big{)}\Big{]}.$ (5)
@@
In the ideal case where there is no optimization error, the minimizer of Equation (5) is the same as the two-staged DRL framework (Rafailov et al., 2023; Azar et al., 2023). Meanwhile, direct preference learning algorithms are generally easier to tune and require fewer computational resources compared to DRL methods. Considering these factors, in our project, we focus on direct preference learning algorithms while leaving the study of the DRL-based framework for future research.
The open-source project Zephyr (Tunstall et al., 2023) serves as a milestone to popularize the DPO algorithm, where the authors provide a comprehensive guide for training LLMs through vanilla DPO and distillation from the teacher model ChatGPT. Following the Zephyr framework, many open-source models have been fine-tuned using vanilla DPO and their qualities are largely improved compared to their SFT counterparts. Impressively, on various LLM benchmarks’ leaderboards, such as those reported by (Dubois et al., 2023; Zheng et al., 2023), models fine-tuned using DPO predominantly exhibit superior alignment and effectiveness.
@@
While the vanilla offline direct preference learning algorithms are useful in some case studies, they also face certain challenges. Specifically, they are considered offline because they learn from an offline preference dataset collected prior to the alignment process. We can formulate this data collection process as:
$\small x\sim d_{0},a^{1}\sim\pi_{D}^{1},a^{2}\sim\pi_{D}^{2},\qquad y\sim\mathrm{Ber}\big{(}\mathbb{P}(a^{1}\succ a^{2}|x,a^{1},a^{2})\big{)}.$ (6)
@@
Here, $(\pi_{D}^{1},\pi_{D}^{2})$ represent two behavior policies, often taken as $\pi_{0}$, other open-sourced models or proprietary models. The term “offline learning” implies that we cannot further query the preference oracle $\mathbb{P}$ during the training process. However, the finite dataset $\mathcal{D}_{\mathrm{off}}$ fails to cover the entire prompt-response space and the resulting policy model often performs poorly when faced with out-of-distribution data (Burns et al., 2023). In particular, along the way of the RLHF training, the average density ratio $\frac{\pi(a|x)}{\pi_{0}(a|x)}>\exp(25)$ as reported in Figure 13 of Bai et al. (2022a). Therefore, the distribution shift between policies is usually very large, and it is unlikely that we can learn the optimal policy solely from a pre-collected dataset.
@@ 1.2 Online Iterative RLHF
In contrast, the Claude project (Bai et al., 2022a) and LLaMA-2 project (Touvron et al., 2023) have demonstrated that online iterative RLHF can significantly improve model performance. The process of online iterative RLHF, as formally formulated in Xiong et al. (2023), can be summarized as follows. Given the pre-collected preference dataset $\mathcal{D}=\mathcal{D}_{\mathrm{off}}$ (if applicable, otherwise empty), for each iteration $t\in[T]$:
@@
The effectiveness of online data can be intuitively understood as continuously collecting new online data to strategically explore the prompt-response space and mitigating the out-of-distribution (OOD) issue. We also refer readers to Xiong et al. (2023) for a detailed theoretical explanation for online iterative RLHF. The hybrid formulation presented here is mainly for generality and is motivated by the LLaMA-2 project (Touvron et al., 2023). This formulation is also loosely related to some of the previous RL literature like Xie et al. (2021); Song et al. (2022).
We notice that the results presented in LLaMA-2 and Claude are based on the deep RL method, PPO, and the data, models, and training details are not fully accessible to the open-source community. Moreover, compared to the vanilla offline DPO, its online iterative version is still largely under-explored in the literature. Xiong et al. (2023) made the first step towards understanding the advantage of online exploration in DPO training from a theoretical perspective, and the main purpose of this work is to provide a detailed guidance to make the online iterative RLHF pipeline more accessible to the open-source community so that others can easily reproduce.
@@ 1.3 Human Feedback Approximation
Ideally, the online preference signal is sampled from a representative group of human labelers. However, human feedback is extremely expensive in practice, which the open-source community usually cannot afford. In the literature, there is a line of work showing that training a proxy preference model, and using the preference model to give proxy labels in a semi-supervised manner improve the model performance (Dong et al., 2023; Yuan et al., 2023; Liu et al., 2023a; Hoang Tran, 2024). We conjecture that this is because the reward model (discriminator) usually generalizes better than the policy (generator).
In particular, Hoang Tran (2024) shows that if the preference model (reward model) is trained on a diverse set of preference datasets, the Pair-RM (Jiang et al., 2023) with only 0.4B parameters can provide iterative preference learning with meaningful signals so that the resulting model [^1] https://huggingface.co/snorkelai/Snorkel-Mistral-PairRM-DPO achieves an impressive AlpacaEval-2 length-control win rate of 26.4%. Motivated by this line of work, we first train a proxy preference (reward) model based on the diverse open-source preference datasets in Section 2 and then use the resulting model to provide preference signals for the subsequent iterative RLHF.
@@ 1.4 Related Work
We have presented many related works in the area of RLHF in the previous subsections. For completeness, we also include a more comprehensive literature review in this subsection.
@@
RLHF and RLHF algorithms. The idea of learning from relative feedback could date back to the study of dueling bandit (Joachims et al., 2007; Yue et al., 2012) and the current RLHF framework was first popularized by Christiano et al. (2017), which served to direct the attention of the deep RL community to the preference-based feedback. Then, these techniques were further introduced to fine-tune LLMs for the summarization task (Ziegler et al., 2019; Stiennon et al., 2020). The dominant RLHF framework used for the modern LLM alignment was first thoroughly developed in Instruct-GPT (ChatGPT) (Ouyang et al., 2022), Claude (Bai et al., 2022a), and LLaMA-2 (Touvron et al., 2023) projects. These works typically involve constructing a reward model based on the MLE of the Bradley-Terry model, and then using the PPO algorithm to optimize the reward signals with KL regularization. One notable exception is that the LLaMA-2 uses a mixture of rejection sampling fine-tuning (Dong et al., 2023; Wang et al., 2024) and PPO in their RLHF pipeline. We refer the interested readers to Bai et al. (2022a); Touvron et al. (2023) for a detailed description. However, the use of PPO in RLHF has limitations. It is known to be unstable (Choshen et al., 2019), sensitive to implementation (Engstrom et al., 2020), and resource-intensive (Yuan et al., 2023). Despite some efforts to improve PPO in the context of RLHF (Li et al., 2023; Chan et al., 2024; Chang et al., 2024; Zhong et al., 2024), reproducing the successful results achieved with PPO is challenging for the open-source community due to these limitations as it requires extensive efforts and resources that the open-source communities usually cannot afford. In recognition of these issues of PPO, a line of work studies the (offline) direct preference learning algorithms, including Slic (Zhao et al., 2023), DPO (Rafailov et al., 2023), IPO (Azar et al., 2023), KTO (Ethayarajh et al., 2024), ARM (Pang et al., 2024), and GPO (Tang et al., 2024). These algorithms skip the reward modeling step, and optimize a designed loss target on the offline preference dataset directly (hence the name). It is widely observed that the direct preference learning algorithms are much more stable than the PPO, and achieve impressive performance evaluated by standard benchmarks (Tunstall et al., 2023; Dubois et al., 2023; Zheng et al., 2023).
Despite the advances offered by these direct preference learning algorithms, their implementation is typically offline and off-policy. This means they operate on preference datasets that were previously collected by other models—often powerful, proprietary teacher models like ChatGPT—before the training process begins (Tunstall et al., 2023; Cui et al., 2023).
@@
RLHF benefits from online (iterative) learning. Roughly speaking, online iterative learning means that we will deploy the intermediate models and query human feedback for the responses of these models. Intuitively, this strategy can help to mitigate the OOD issue of the learned reward model (Gao et al., 2023), and its advantages have been reported in Ouyang et al. (2022); Touvron et al. (2023) for the PPO-based framework. Even when the additional feedback is derived from a proxy reward constructed from the same offline dataset (similar to semi-supervised learning), iterative rejection sampling fine-tuning (RAFT) (Dong et al., 2023) and DPO based on samples from the target distribution estimator have been shown to outperform the original offline counterparts (Pang et al., 2024; Liu et al., 2023a). Furthermore, recent works (Xiong et al., 2023; Xu et al., 2023b; Hoang Tran, 2024; Yuan et al., 2024b; Swamy et al., 2024; Chen et al., 2024b; Ye et al., 2024; Guo et al., 2024; Rosset et al., 2024; Tajwar et al., 2024; Calandriello et al., 2024; Wu et al., 2024) have demonstrated that online iterative variants of direct preference learning algorithms significantly outperform their offline counterparts. In particular, we refer the interested readers to Guo et al. (2024) for the extensive experimental results with different offline base algorithms.